copper react with oxygen to form two oxide x and y analysis 1.535g of x yield 1.365g of copper and 1.450g of y yield 1.60g of copper . determine the chemical formula x and y, calculate the mass of copper which can react with 0.500g of oxygen to form x and y


1
Expert's answer
2021-03-11T08:46:50-0500

1) 1.535 g of X -> 1.365 g of Cu

1.535 – 1.365 = 0.170g of O

Aw of Cu =63.5, of O = 16

For Cu 1.365 g / 63.5 = 0.02 mol

For O 0.170 g / 16 = 0.01 mol

X = Cu2O

There is missing information about mass of Y yielded of copper, which is necessary for solving this task, so I tried to find the full question and I will be giving answer according to the founded missing part.

Copper reacts with oxygen to form two oxides X and Y on analysis, 1.535g of X yielded 1.365g if copper and 1.450g of Y yield 1.160g of copper. Determine the chemical formula of X and Y and calculate the mass of copper which can react with 0.500g of oxygen to yield X and Y?


2) 1.450 g of Y -> 1.160 g of Cu

1.450 – 1.160 = 0.290 g of O

For Cu = 1.160 g / 63.5 = 0.018 mol

For O = 0.290 g / 16 = 0.018 mol

Y = CuO


3) total mass of O = 0,170 g + 0,290 g = 0.460 g

Total mass of Cu = 1.160 g + 1. 365 g = 2.525 g

0.460 g of O -> 2.525 g of Cu

0.500 g of O -> (2.525 x 0.5) / 0.460 = 2.745 g of Cu